THE ELECTRONIC STRUCTURE OF ScH^{++}, TiH^{++}, VH^{++}, CrH^{++},AND MnH^{++}

\underline{JAMES F. HARRISON}, PETER S. CHRISTOPHER, Chemistry Department and Center For Fundamental Materials Research, Michigan State University, East Lansing, MI 48824; , .

We will discuss the electronic structure of the ground and low-lying excited states of the titled transition metal hydrides. Of these ScH^{++}, TiH^{++},and VH^{++}, are thermodynamically stable while CrH^{++}, and MnH^{++} are expected to be very long lived. While the charge distribution for each molecule is approximately M^{+1.85} H^{+0.15} the nature of the bonding changes very rapidly from a conventional sigma bond in ScH^{++}, to what may be described as an antiferromagnetic coupling of the Mn^{++} ion to the H atom in MnH^{++}. We will discuss trends in D_{e} , bond lengths and vibrational frequencies and will compare with the corresponding neutral and monopositive hydrides.